OpenStudy (anonymous):

A city has a population of 284,900 which is 110% of the population last year. What was the population last year? can someone confirm my answer A:The population was 257,000 last year.

OpenStudy (ash2326):

110x/100=284900 110x=284900*100 x=259000 you recheck your solution